# Distillation of Discrete Diffusion through Dimensional Correlations

### v3

@@ Acknowledgments
The authors would like to thank Chieh-Hsin Lai, Naoki Murata, Bac Nguyen, and Toshimitsu Uesaka for their constructive feedback.
@@ A.4 Iterated Di4C training
While we generally assume that the teacher model is given by a product model, the Di4C loss functions can also treat mixture teacher models. The only exception is the marginal loss (Section A.3), but we can just replace $q_{0|t}$ with $\overline{q}_{0|t}$ in (14) and conduct Monte Carlo estimates. Through this generalization, we can run multiple rounds of Di4C in a similar spirit as the multi-round SDTT (Deschenaux & Gulcehre, 2024).
@@ D.5 Proof of Proposition 6
$p_{s|t}(aa|aa)+p_{s|t}(bb|aa)$ $=\left(\frac{(1+e^{-(t+s)})(1+e^{-(t-s)})}{2(1+e^{-2t})}\right)^{2}+\left(\frac{(1-e^{-(t+s)})(1-e^{-(t-s)})}{2(1+e^{-2t})}\right)^{2}$
@@
$\frac{e^{-2\epsilon}-e^{-2(t-\epsilon)}}{1-e^{-2t}}-\frac{(e^{-\epsilon}-e^{-(2t-\epsilon)})^{2}}{(1-e^{-2t})^{2}}$ $=\frac{(e^{-2\epsilon}-e^{-2(t-\epsilon)})(1-e^{-2t})-(e^{-\epsilon}-e^{-(2t-\epsilon)})^{2}}{(1-e^{-2t})^{2}}$
@@
$\frac{e^{-2\epsilon}+e^{-2(t-\epsilon)}}{1+e^{-2t}}-\frac{(e^{-\epsilon}+e^{-(2t-\epsilon)})^{2}}{(1+e^{-2t})^{2}}$ $=\frac{(e^{-2\epsilon}+e^{-2(t-\epsilon)})(1+e^{-2t})-(e^{-\epsilon}+e^{-(2t-\epsilon)})^{2}}{(1+e^{-2t})^{2}}$
@@
$\frac{e^{-2\epsilon}+e^{-2(t-\epsilon)}}{1+e^{-2t}}+\frac{e^{-2\epsilon}-e^{-2(t-\epsilon)}}{1-e^{-2t}}$ $=\frac{(e^{-2\epsilon}+e^{-2(t-\epsilon)})(1-e^{-2t})+(e^{-2\epsilon}-e^{-2(t-\epsilon)})(1+e^{-2t})}{1-e^{-4t}}$
@@
$\frac{e^{-2t}}{2(1+e^{-2t})^{2}}p_{t}^{\epsilon}(aa)+\frac{e^{-2t}}{2(1-e^{-2t})^{2}}\left(\frac{1}{2}-p_{t}^{\epsilon}(aa)\right)$ $\geq\frac{1}{2}\min\left\{\frac{e^{-2t}}{2(1+e^{-2t})^{2}},\ \frac{e^{-2t}}{2(1-e^{-2t})^{2}}\right\}=\frac{e^{-2t}}{4(1-e^{-2t})^{2}}.$
@@
$\left(\frac{e^{-2t}}{2(1+e^{-2t})^{2}}p_{t}^{\epsilon}(aa)+\frac{e^{-2t}}{2(1-e^{-2t})^{2}}\left(\frac{1}{2}-p_{t}^{\epsilon}(aa)\right)\right)(e^{\epsilon}-e^{-\epsilon})^{2}$ $\geq\frac{e^{-2t}}{4(1-e^{-2t})^{2}}\cdot 4\epsilon^{2}=\frac{e^{-2t}}{(1-e^{-2t})^{2}}\epsilon^{2}.$ (58)
@@ E.2 Derivations of dimension-wise computable control variates for mixture model
$\mathbb{E}_{\bm{x}\sim q}\mathbb{E}_{\lambda}\!\left[-\log p^{\theta,\lambda}(\bm{x})\right]$ $=\mathbb{E}_{\eta}\mathbb{E}_{\bm{x}\sim q^{\eta}}\mathbb{E}_{\lambda}\!\left[-\log p^{\theta,\lambda}(\bm{x})\right]=\mathbb{E}_{\eta}\mathbb{E}_{\lambda}\mathbb{E}_{\bm{x}\sim q^{\eta}}\!\left[-\log p^{\theta,\lambda}(\bm{x})\right]$
@@ Appendix F Experimental details
F.1 Latency overhead of mixture modeling
Experiment # steps teacher (sec.) student (sec.)
CIFAR-10 10 0.5515$\pm$0.0024 0.5786$\pm$0.0017
ImageNet 4 2.0741$\pm$0.0035 2.0734$\pm$0.0043
OpenWebText 16 3.3409$\pm$0.0417 3.4817$\pm$0.0730
F.2 Discretized Gaussian diffusions
F.2.1 Sampling schemes
@@
F.2.2 Detailed experimental results
@@ F.2.5 Generated samples
F.3 Masked generative image modeling
F.3.1 Masked diffusion modeling
@@
F.3.2 Confidence-based sampling
@@
F.3.3 Discrete classifier-free guidance
@@ F.3.5 Generated samples
Samples generated by each model in Section 5.2 are shown in Figure 6. The best CFG coefficient in Figure 3 in terms of FID was chosen for each model, as shown in Table 6 together with the FID/IS performance. Regarding the choice of ImageNet labels (Ostrich, Burger, Volcano), we followed Besnier & Chen (2023, Figure 3).
@@
Model # steps $w_{\mathrm{cfg}}$ FID IS
teacher 8 3 6.57 202.0
teacher 4 11 7.97 216.0
di4c 4 6 6.79 209.2
di4c-d 4 7 6.57 213.6
F.4 Masked diffusion language models
F.4.1 Diffusion modeling and sampling
@@ F.4.2 Implementation and training
For the teacher model, we used the two “small" checkpoints (round $6$ and $7$) of Deschenaux & Gulcehre (2024)333 It is loaded by load_small_student(loss='kld', round=$n$) with $n=6,7$, from the library sdtt in https://github.com/jdeschena/sdtt. , which uses a transformer architecture to compute the logits of $p_{0|t}(\cdot|\bm{x}_{t})$ for each token. The transformer architecture is originally from (Sahoo et al., 2024), and has 169M parameters with $12$ layers, the embedding dimension of $768$, and $12$ attention heads. The model receives $D=1024$ tokens and does not depend on the timestep. For adaptation to mixture modeling, we just applied the same modification as given in Section F.3.4.
@@
Except for the Self-BLEU computation, all the evaluations were done using the code of SDTT (Deschenaux & Gulcehre, 2024) [^6] https://github.com/jdeschena/sdtt/tree/main..
F.4.3 Additional experimental results
@@
We also tested our models with the MAUVE score (Pillutla et al., 2021). The setting is the same as the unconditional generation in Section 5.3, and the MAUVE computation is done by using the code of SDTT (Deschenaux & Gulcehre, 2024). As shown in Figure 8, no significant performance decay from the teacher model was observed.

### v4

@@ 7 Conclusion
Impact Statement
This paper presents a novel method for reducing the number of sampling steps required for high-quality generation in discrete data domains. While there are many potential social impacts as usual in the generative AI field, we do not foresee any specific problems particularly caused by this (mostly theoretical) work.
@@ Acknowledgments
We would like to thank Masato Ishii, Chieh-Hsin Lai, Naoki Murata, Bac Nguyen, and Toshimitsu Uesaka for their constructive feedback. We also thank the anonymous reviewers for their valuable comments and suggestions.
@@ A.4 Iterated Di4C training
While we generally assume that the teacher model is given by a product model, the Di4C loss functions can also treat mixture teacher models. The only exception is the marginal loss (Section A.3), but we can just replace $q_{0|t}$ with $\overline{q}_{0|t}$ in (14) and conduct Monte Carlo estimates. Through this generalization, we can run multiple rounds of Di4C in a similar spirit as the multi-round SDTT (Deschenaux & Gulcehre, 2025).
@@ D.5 Proof of Proposition 6
$p_{s|t}(aa|aa)+p_{s|t}(bb|aa)$
$=\left(\frac{(1+e^{-(t+s)})(1+e^{-(t-s)})}{2(1+e^{-2t})}\right)^{2}+\left(\frac{(1-e^{-(t+s)})(1-e^{-(t-s)})}{2(1+e^{-2t})}\right)^{2}$
@@
$\frac{e^{-2\epsilon}-e^{-2(t-\epsilon)}}{1-e^{-2t}}-\frac{(e^{-\epsilon}-e^{-(2t-\epsilon)})^{2}}{(1-e^{-2t})^{2}}$
$=\frac{(e^{-2\epsilon}-e^{-2(t-\epsilon)})(1-e^{-2t})-(e^{-\epsilon}-e^{-(2t-\epsilon)})^{2}}{(1-e^{-2t})^{2}}$
@@
$\frac{e^{-2\epsilon}+e^{-2(t-\epsilon)}}{1+e^{-2t}}-\frac{(e^{-\epsilon}+e^{-(2t-\epsilon)})^{2}}{(1+e^{-2t})^{2}}$
$=\frac{(e^{-2\epsilon}+e^{-2(t-\epsilon)})(1+e^{-2t})-(e^{-\epsilon}+e^{-(2t-\epsilon)})^{2}}{(1+e^{-2t})^{2}}$
@@
$\frac{e^{-2\epsilon}+e^{-2(t-\epsilon)}}{1+e^{-2t}}+\frac{e^{-2\epsilon}-e^{-2(t-\epsilon)}}{1-e^{-2t}}$
$=\frac{(e^{-2\epsilon}+e^{-2(t-\epsilon)})(1-e^{-2t})+(e^{-2\epsilon}-e^{-2(t-\epsilon)})(1+e^{-2t})}{1-e^{-4t}}$
@@
$\frac{e^{-2t}}{2(1+e^{-2t})^{2}}p_{t}^{\epsilon}(aa)+\frac{e^{-2t}}{2(1-e^{-2t})^{2}}\left(\frac{1}{2}-p_{t}^{\epsilon}(aa)\right)$
$\geq\frac{1}{2}\min\left\{\frac{e^{-2t}}{2(1+e^{-2t})^{2}},\ \frac{e^{-2t}}{2(1-e^{-2t})^{2}}\right\}=\frac{e^{-2t}}{4(1-e^{-2t})^{2}}.$
@@
$\left(\frac{e^{-2t}}{2(1+e^{-2t})^{2}}p_{t}^{\epsilon}(aa)+\frac{e^{-2t}}{2(1-e^{-2t})^{2}}\left(\frac{1}{2}-p_{t}^{\epsilon}(aa)\right)\right)(e^{\epsilon}-e^{-\epsilon})^{2}$
$\geq\frac{e^{-2t}}{4(1-e^{-2t})^{2}}\cdot 4\epsilon^{2}=\frac{e^{-2t}}{(1-e^{-2t})^{2}}\epsilon^{2}.$ (58)
@@ E.2 Derivations of dimension-wise computable control variates for mixture model
$\mathbb{E}_{\bm{x}\sim q}\mathbb{E}_{\lambda}\!\left[-\log p^{\theta,\lambda}(\bm{x})\right]$
$=\mathbb{E}_{\eta}\mathbb{E}_{\bm{x}\sim q^{\eta}}\mathbb{E}_{\lambda}\!\left[-\log p^{\theta,\lambda}(\bm{x})\right]=\mathbb{E}_{\eta}\mathbb{E}_{\lambda}\mathbb{E}_{\bm{x}\sim q^{\eta}}\!\left[-\log p^{\theta,\lambda}(\bm{x})\right]$
@@ Appendix F Experimental details
F.1 Discretized Gaussian diffusions
F.1.1 Sampling schemes
@@
F.1.2 Additional experimental results
@@
FID/IS results of fewer sampling steps.
For additional comparison, we computed FID/IS of the teacher and student modles using 10K samples (fewer than the 50K samples used in the main body, so the numbers may be slightly worse) for 2-20 steps in Table 4. In terms of FID, our method achieves approximately 1.4 times acceleration in the 10–20 steps range. However, it does not perform well in very few steps (e.g., 2–4 steps).
Table 4: Comparison of models on CIFAR-10 dataset in various sampling steps. Same setting as Table 1 except that 10,000 generated samples were used for computing FID/IS.
#steps 2 4 6 8 10 12 14 16 18 20
FID teacher 392.24 173.29 78.24 49.44 34.70 26.33 21.47 18.05 15.80 14.42
student 411.70 147.67 59.62 33.85 22.57 17.46 14.37 12.86 12.28 11.81
IS teacher 1.17 2.99 5.90 7.01 7.45 7.82 7.96 8.16 8.35 8.46
student 1.25 3.48 6.71 7.68 8.17 8.33 8.37 8.50 8.38 8.39
More detailed results of Table 1.
@@ F.1.5 Generated samples
F.2 Masked generative image modeling
F.2.1 Masked diffusion modeling
@@
F.2.2 Confidence-based sampling
@@
F.2.3 Discrete classifier-free guidance
@@ F.2.5 Generated samples
Samples generated by each model in Section 5.2 are shown in Figure 6. The best CFG coefficient in Figure 3 in terms of FID was chosen for each model, as shown in Table 8. The table also shows the FID/IS performance, together with the Precision/Recall values computed by the code of Besnier & Chen (2023). Regarding the choice of ImageNet labels (Ostrich, Burger, Volcano), we followed Besnier & Chen (2023, Figure 3).
@@
Model # steps $w_{\mathrm{cfg}}$ FID ($\downarrow$) IS ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$)
teacher 8 3.0 6.57 202.0 0.7939 0.5499
teacher 4 11.0 7.97 216.0 0.7737 0.5057
di4c 4 6.0 6.79 209.2 0.7910 0.5363
di4c-d 4 7.0 6.57 213.6 0.7866 0.5391
F.3 Masked diffusion language models
F.3.1 Diffusion modeling and sampling
@@ F.3.2 Implementation and training
For the teacher model, we used the two “small” checkpoints (round $6$ and $7$) of Deschenaux & Gulcehre (2025)333 It is loaded by load_small_student(loss='kld', round=$n$) with $n=6,7$, from the library sdtt in https://github.com/jdeschena/sdtt. , which uses a transformer architecture to compute the logits of $p_{0|t}(\cdot|\bm{x}_{t})$ for each token. The transformer architecture is originally from (Sahoo et al., 2024), and has 169M parameters with $12$ layers, the embedding dimension of $768$, and $12$ attention heads. The model receives $D=1024$ tokens and does not depend on the timestep. For adaptation to mixture modeling, we just applied the same modification as given in Section F.2.4.
@@
Except for the Self-BLEU computation, all the evaluations were done using the code of SDTT (Deschenaux & Gulcehre, 2025) [^6] https://github.com/jdeschena/sdtt/tree/main..
F.3.3 Additional experimental results
@@
We also tested our models with the MAUVE score (Pillutla et al., 2021). The setting is the same as the unconditional generation in Section 5.3, and the MAUVE computation is done by using the code of SDTT (Deschenaux & Gulcehre, 2025). As shown in Figure 8, no significant performance decay from the teacher model was observed.